# Simulation Based Bayesian Optimization

### v1

@@ 3.1 Main Idea
If the utility is non-negative and bounded and the state space of the inhomogeneous Markov chain produced by this sampler is discrete, then Müller et al. (2004) show that the chain is strongly ergodic if the updating schedule for $H$ is logarithmic. Moreover, its stationary distribution is uniform over the set of $\boldsymbol{x}$ maximizing (3.1).
A key point of this approach is that once a candidate $\tilde{\boldsymbol{x}}_{n+1}$ has been proposed, the distribution $\pi(f(\boldsymbol{x})|\tilde{\boldsymbol{x}}_{n+1},\mathcal{D}_{1:n})$ is used as proposal for generating $\tilde{f}(\boldsymbol{x})_{h}$ for $h=1,\dots,H$. As a consequence, within each iteration of Algorithm 1, the acceptance ratio is:
$\alpha$ $=$ $\min\left\{1,\frac{\prod_{h=1}^{H}u(\tilde{f}(\boldsymbol{x})_{h},\tilde{\boldsymbol{x}}_{n+1})\pi(\tilde{f}(\boldsymbol{x})_{h}|\tilde{\boldsymbol{x}}_{n+1},\mathcal{D}_{1:n})\pi(f(\boldsymbol{x})^{(0)}_{h}|\boldsymbol{x}^{(0)}_{n+1},\mathcal{D}_{1:n})q(\boldsymbol{x}^{(0)}_{n+1}|\tilde{\boldsymbol{x}}_{n+1})}{\prod_{h=1}^{H}u(f(\boldsymbol{x})^{(0)}_{h},\boldsymbol{x}^{(0)}_{n+1})\pi(f(\boldsymbol{x})^{(0)}_{h}|\boldsymbol{x}^{(0)}_{n+1},\mathcal{D}_{1:n})\pi(\tilde{f}(\boldsymbol{x})_{h}|\tilde{\boldsymbol{x}}_{n+1},\mathcal{D}_{1:n})q(\tilde{\boldsymbol{x}}_{n+1}|\boldsymbol{x}^{(0)}_{n+1})}\right\}$
$=$ $\min\left\{1,\frac{\prod_{h=1}^{H}u(\tilde{f}(\boldsymbol{x})_{h},\tilde{\boldsymbol{x}}_{n+1})q(\boldsymbol{x}^{(0)}_{n+1}|\tilde{\boldsymbol{x}}_{n+1})}{\prod_{h=1}^{H}u(f(\boldsymbol{x})^{(0)}_{h},\boldsymbol{x}^{(0)}_{n+1})q(\tilde{\boldsymbol{x}}_{n+1}|\boldsymbol{x}^{(0)}_{n+1})}\right\}$
$=$ $\min\left\{1,\exp(H\tilde{v}-Hv^{(0)})\frac{q(\boldsymbol{x}^{(0)}_{n+1}|\tilde{\boldsymbol{x}}_{n+1})}{q(\tilde{\boldsymbol{x}}_{n+1}|\boldsymbol{x}^{(0)}_{n+1})}\right\}.$
Thus, the posterior predictive density for $\tilde{f}(\boldsymbol{x})$ cancels out with the proposal density and we do not need to evaluate it. This allows us to optimize (3.1) when just having sampling access to the posterior, thus achieving our goal.
@@ 3.2 Gibbs Sampler
$\begin{split}&g_{H}(\boldsymbol{x}_{s}|\boldsymbol{x}_{-s},f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H}),\mathcal{D}_{1:n})\propto\\ &\quad\exp\Bigg{\{}\sum_{h=1}^{H}\log\left[u(f(\boldsymbol{x})_{h},\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s}\right]+\log\left[\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s},\mathcal{D}_{1:n})\right]\Bigg{\}}\end{split}$ (3.5)
For the particular case in which $\boldsymbol{x}$ is a vector of $p$ categorical variables, this marginal is simply the softmax distribution over $\sum_{h=1}^{H}\log\left[u(f(\boldsymbol{x})_{h},\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s}\right]+\log\left[\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s},\mathcal{D}_{1:n})\right]$. However, sampling from this distribution requires evaluating posterior predictive probabilities, which defeats the purpose of SBBO. As an alternative, we could create a Gibbs sampling scheme to sample from $g_{H}(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{H}|\mathcal{D}_{1:n})$ where $(f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H})$ has been integrated out. Sampling from $g_{H}(\boldsymbol{x}_{s}|\boldsymbol{x}_{-s})$ can be done using Metropolis steps. Assuming that the current state of the chain is $\boldsymbol{x}$, $f(\boldsymbol{x})_{1},\dots,f(\boldsymbol{x})_{H}$ and $v=\frac{1}{H}\sum{h=1}^{H}\log u(f(\boldsymbol{x})_{h},\boldsymbol{x})$, we can proceed as follows:
@@ 3.3 Probabilistic Surrogate Models
Tanimoto Gaussian Process.
This surrogate model is particularly useful when $x\in\{0,1\}^{p}$. The uncertainty on $f(x)$ is modeled through a Gaussian Process with the following kernel function:
$k(x,x^{\prime})=\frac{\phi(x\cdot x^{\prime})}{\|x\|^{2}+\|x^{\prime}\|^{2}-x\cdot x^{\prime}}.$
The numerator is proportional to the number of dimensions where both $x$ and $x^{\prime}$ are 1, and the denominator signifies the total number of dimensions where either $x$ or $x^{\prime}$ or both is 1. The parameter $\phi$ is usually set to the value maximizing the marginal likelihood thus following a empirical Bayes approach. For categorical covariates with more than two possible values, we would need to first convert the covariates into appropriate binary variables (e.g. via one-hot encoding) before computing the previous kernel. Under this model, the posterior predictive distribution is Gaussian with a certain mean and variance analytically available as in any GP.
Sparse Bayesian linear regression.
@@
$\epsilon\sim\mathcal{N}(0,\sigma^{2})$ (3.7)
@@
This model was first proposed by Baptista and Poloczek (2018) in the context of Bayesian Optimization. Posterior inference is achieved through Gibbs sampling, as detailed in Baptista and Poloczek (2018). This method allows us to sample from the posterior of $\alpha_{k},\beta_{k},\tau$, and $\sigma^{2}$, facilitating easy draws from the posterior of $f_{\alpha}(\boldsymbol{x})$ that could be directly utilized in Algorithm 1. In this scenario, access to analytical posterior predictive distributions is unavailable, making the computation of common acquisition functions challenging. To address this, Baptista and Poloczek (2018) propose a tailored acquisition function: a single posterior sample of $\alpha$ is produced, and the acquisition function is defined as $f_{\alpha}(\boldsymbol{x})-\lambda\|\boldsymbol{x}\|_{1}$, where a penalty term controlled by $\lambda$ is included. This acquisition function now has a closed-form, appearing as a quadratic form. Its optimization can be reformulated as a semidefinite program that can be efficiently approximated in polynomial time to a desired precision. With SBBO, there is no need for tailored acquisition functions as we just require having sampling access to the posterior predictive distribution. This enables using model (3.6) with the common acquistion functions in the BO literature.
NGBoost.
@@
Natural Gradient Boosting (NGBoost, Duan et al. (2020)) is a non-Bayesian algorithm for generic probabilistic prediction via gradient boosting. In NGBoost, the conditional distribution of the output given covariates is modelled through a parametric distribution $y|\boldsymbol{x}\sim P_{\theta}(\boldsymbol{x})$. The parameters $\theta(\boldsymbol{x})$ are obtained through an additive combination of $M$ base learners and an initial $\theta^{(0)}$.
$\theta=\theta^{(0)}-\eta\sum_{m=1}^{M}\rho^{(m)}\cdot f^{(m)}(x)$
@@
where $\rho^{(m)}$ are scaling factors and $\eta$ is a common learning rate. The base learners are trained to minimize a proper scoring rule using a refinement of the gradient boosting algorithm. For this training process, NGBoost employs the natural gradient, a variant of gradient descent that takes into account the underlying geometry of probability distributions. This leads to faster convergence and improved stability during training.
@@
A notable strength of NGBoost lies in its adaptability, as it can be paired with various base learners. Our experiments focus on two such learners: shallow decision trees (referred to as NGBoost-dec) and linear regressions with lasso regularization (named NGBoost-linCV). For both these configurations, we utilize the log-score as the scoring rule.
Bayesian Neural Network.
@@
Bayesian Neural Networks (BNNs) are probabilistic deep learning models that replace deterministic weight values of a neural network with weight distributions. This allows the model to capture the existing epistemic uncertainty in the predictions due to limited training data. In our experiments, we employ a simple three- hidden layers fully-connected BNN whose inputs are the one-hot encoded covariates. We train this model using variational inference, combined with the reparametrization estimator technique (Kingma and Welling, 2013).
@@ 4 Experiments
We conduct experiments in three different combinatorial optimization problems [^1] All code to reproduce the experiments in this paper has been open sourced and is available at https://github.com/roinaveiro/sbbo.: a binary quadratic program with 10 variables, a contamination control in a food supply chain with 25 stages, and RNA sequence optimization with a sequence length of 30. Details of the problem settings are provided in the subsequent subsections. We evaluate the performance of SBBO with the various probabilistic surrogate models described at the end of the previous section: Sparse Bayesian linear regression with interactions (sbbo-BLr), Bayesian Neural Network (sbbo-BNN), Tanimoto Gaussian Process Regression (sbbo-GPr), natural gradient boosting with a shallow decision tree model as base learner (sbbo-NGBdec), and natural gradient boosting with an sparse linear regression as base learner (sbbo-NGBlinCV).
We compare performance of these SBBO approaches with two benchmarks: simulated annealing (SA), a widely used algorithm in combinatorial optimization problems (Spears, 1993) and random local search (RS, Bergstra and Bengio (2012)).
The experimental setup was design as follows. We start with a random initial dataset containing five pairs of covariates and responses. For each probabilistic surrogate model and fixed number of iterations $t$, we follow these steps:
@@
    Use the surrogate model to determine the next evaluation point, utilizing the expected improvement as the acquisition function. Thus, the utility function is defined as $u(\boldsymbol{x})=\max\left[f(\boldsymbol{x})-f^{*}(\boldsymbol{x}),0\right]$, where $f^{*}(\boldsymbol{x})$ represents the current estimate of the optimal true objective. For SBBO algorithms, we employ a Metropolis-Hastings sampling scheme, as described in Algorithm 1. The proposal distribution randomly selects one of the covariates and proposes another value, chosen uniformly at random, which is accepted according to the acceptance probability in step 4 of Algorithm 1. The chosen cooling schedule incrementally increases $H$ from 1 to 10000 in steps of 250.
@@
For each algorithm, we report the best function value found after $t$ evaluations of the true objective, averaged over 10 runs, with one standard deviation divided by the square root of the number of runs as the margin of error. Note that $t$ varies depending on the problem setting.
@@ 4.1 Binary Quadratic Problem
Table 1: Average objective function values and margin of error for each algorithm, obtained after 120 iterations for the binary quadratic problem, 500 iterations for the contamination control problem and 300 iterations for the RNA problem. Bold values denote the best value for each problem.
@@
The aim here is to maximize a quadratic function while incorporating $l_{1}$ regularization. This objective function can be expressed as $\boldsymbol{x}^{\top}Q\boldsymbol{x}-\lambda\|\boldsymbol{x}\|_{1}$, where $\boldsymbol{x}$ is a binary vector of dimension $d=10$. The matrix $Q$ is a random $d\times d$ matrix with entries generated as follows: each element is simulated as independent standard Gaussian, and the resulting matrix is then element-wise multiplied by another $d\times d$ matrix $K$. The elements of matrix $K$ are defined as $K_{ij}=\exp\left[-(i-j)^{2}/L_{c}^{2}\right]$, where the correlation length $L_{c}^{2}$ governs how rapidly the values of $K$ decay away from the diagonal. A larger value of $L_{c}^{2}$ results in a denser matrix $Q$, which makes the optimization task more challenging. In our experiments, we set $\lambda$ to 0 and fixed $L_{c}^{2}$ at 10.
Figure 1: BQP problem: average objective function values by number of evaluations, colored by algorithm. Large objective values indicate better performance.
The performance comparison of the different algorithms over 120 iterations is illustrated in Figure 1 and detailed in the second column of Table 1. As can be seen, SBBO outperforms RS and SA for most surrogate models. Notably, SBBO with sparse Bayesian linear regression achieves the highest objective value, closely followed by SBBO with NGB employing a sparse linear model as the base learner. An essential aspect is the assessment of performance with a limited number of objective function evaluations, as evaluating the true objective is the common bottleneck in most BO applications. Remarkably, SBBO-BLr, SBBO-BNN, and SA demonstrate good performance for low number of evaluations ($<$25) (Fig. 1). However, it is worth noting that SA appears to get stuck in local optima around the 65th evaluation and SBBO-BNN performance does not grow as much as that of SBBO-BLr.
In this experiment, with a search space cardinality of 1024, it is feasible to compute the global maximizer of the objective function via complete ennumeration (11.24). Table 2 displays the deviation between the optimal solution discovered within 120 function evaluations for each algorithm and the global optimum. Notably, sbbo-BLr consistently identifies the optimal solution within 120 $\ll$ 1024 iterations across all 10 repetitions of the experiment.
Table 2: Binary quadratic problem: average distance to global optimum achieved after 120 iterations with margin of error. Bold indicates best average performance.

### v2

@@ 3.1 Main Idea
If the utility is non-negative and bounded, and the state space of the inhomogeneous Markov chain generated by the sampler is finite, [24] demonstrates that the chain is strongly ergodic when the update schedule for $H$ is logarithmic. Additionally, its stationary distribution is uniform over the set of $\boldsymbol{x}$ that maximize (2.4). In practice, when the state space is not finite, it will be necessary to discretize $v$ into a finite grid.
We have now transformed the optimization of (2.4) into a simulation problem. The objective is to solve this optimization without explicitly evaluating the posterior $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$, relying only on samples from it. Therefore, we need to simulate from the augmented distribution (3.1) without requiring an analytical expression for the posterior. This is where a key observation becomes crucial. Within the MCMC in Algorithm (1), we propose using $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$ as the proposal distribution for $\tilde{f}(\boldsymbol{x})_{h}$ for $h=1,\dots,H$, once a candidate $\tilde{\boldsymbol{x}}_{n+1}$ has been proposed. This approach ensures that the acceptance probability becomes independent of the posterior density, allowing us to simulate from (3.1) without evaluating the posterior directly. To see this more clearly: if the current state of the chain is $(\boldsymbol{x},v)$, where $v=\frac{1}{H}\sum_{h=1}^{H}\log u(f(\mathbf{x})_{h},\mathbf{x})$, the proposed new state is $(\tilde{\boldsymbol{x}},\tilde{v})$, where $\tilde{v}=\frac{1}{H}\sum_{h=1}^{H}\log u(\tilde{f}(\mathbf{x})_{h},\tilde{\boldsymbol{x}})$. With $q(\cdot|\boldsymbol{x})$ as the proposal distribution for $\boldsymbol{x}$, the acceptance ratio for each iteration of Algorithm 1 is:
$\alpha$ $=$ $\min\left\{1,\frac{\prod_{h=1}^{H}u(\tilde{f}(\boldsymbol{x})_{h},\tilde{\boldsymbol{x}})\pi(\tilde{f}(\boldsymbol{x})_{h}|\tilde{\boldsymbol{x}},\mathcal{D}_{1:n})\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x},\mathcal{D}_{1:n})}{\prod_{h=1}^{H}u(f(\boldsymbol{x})_{h},\boldsymbol{x})\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x},\mathcal{D}_{1:n})\pi(\tilde{f}(\boldsymbol{x})_{h}|\tilde{\boldsymbol{x}},\mathcal{D}_{1:n})}\cdot\frac{q(\boldsymbol{x}|\tilde{\boldsymbol{x}})}{q(\tilde{\boldsymbol{x}}|\boldsymbol{x})}\right\}$
$=$ $\min\left\{1,\frac{\prod_{h=1}^{H}u(\tilde{f}(\boldsymbol{x})_{h},\tilde{\boldsymbol{x}})}{\prod_{h=1}^{H}u(f(\boldsymbol{x})_{h},\boldsymbol{x})}\cdot\frac{q(\boldsymbol{x}|\tilde{\boldsymbol{x}})}{q(\tilde{\boldsymbol{x}}|\boldsymbol{x})}\right\}$
$=$ $\min\left\{1,\exp(H\tilde{v}-Hv)\cdot\frac{q(\boldsymbol{x}|\tilde{\boldsymbol{x}})}{q(\tilde{\boldsymbol{x}}|\boldsymbol{x})}\right\}.$
Thus, the posterior predictive density for $\tilde{f}(\boldsymbol{x})$ cancels out with the proposal density, eliminating the need to evaluate it. This allows us to optimize (2.4) when just having sampling access to the posterior, thus achieving our goal.
@@ 3.2 Gibbs Sampler
$\begin{split}&g_{H}(\boldsymbol{x}_{s}|\boldsymbol{x}_{-s},f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H}),\mathcal{D}_{1:n})\propto\\ &\quad\exp\Bigg{\{}\sum_{h=1}^{H}\log\left[u(f(\boldsymbol{x})_{h},\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s})\right]+\log\left[\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s},\mathcal{D}_{1:n})\right]\Bigg{\}}\end{split}$
For the particular case in which $\boldsymbol{x}$ is a vector of $p$ categorical variables, this marginal is simply the softmax distribution over $\sum_{h=1}^{H}\log\left[u(f(\boldsymbol{x})_{h},\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s})\right]+\log\left[\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s},\mathcal{D}_{1:n})\right]$. However, sampling from this distribution requires evaluating posterior predictive probabilities, which defeats the purpose of SBBO. As an alternative, we could create a Gibbs sampling scheme to sample from $g_{H}(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{H}|\mathcal{D}_{1:n})$ where $(f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H})$ has been integrated out. Sampling from $g_{H}(\boldsymbol{x}_{s}|\boldsymbol{x}_{-s},,\mathcal{D}_{1:n})$ can be done using Metropolis steps. Assuming that the current state of the chain is $\boldsymbol{x}$, $f(\boldsymbol{x})_{1},\dots,f(\boldsymbol{x})_{H}$ and $v=\frac{1}{H}\sum_{h=1}^{H}\log u(f(\boldsymbol{x})_{h},\boldsymbol{x})$, we proceed as follows:
@@ 3.3 Probabilistic Surrogate Models
Tanimoto Gaussian Process
This surrogate model, introduced by [22], is particularly useful when $x\in\{0,1\}^{p}$. The uncertainty on $f(x)$ is modeled through a Gaussian Process with the following kernel function:
$k(x,x^{\prime})=\frac{\phi\cdot x^{\top}\cdot x^{\prime}}{\|x\|^{2}+\|x^{\prime}\|^{2}-x\cdot x^{\prime}}.$
The numerator is proportional to the number of dimensions where both $x$ and $x^{\prime}$ are 1, and the denominator is the total number of dimensions where either $x$ or $x^{\prime}$ or both is 1. The parameter $\phi$ is usually set to the value maximizing the marginal likelihood, thus following a empirical Bayes approach. For categorical covariates with more than two possible values, we would need to first convert the covariates into appropriate binary variables (e.g., via one-hot encoding) before computing the previous kernel. Under this model, the posterior predictive distribution is Gaussian with mean and variance analytically available as in any GP.
Sparse Bayesian linear regression
@@
$\epsilon\sim\mathcal{N}(0,\sigma^{2})$
@@
This model was first proposed by [2] in the context of Bayesian Optimization, where posterior inference is performed using Gibbs sampling. This method allows us to sample from the posterior of $\alpha_{k},\beta_{k},\tau$, and $\sigma^{2}$, facilitating easy draws from the posterior of $f_{\alpha}(\boldsymbol{x})$ that could be directly utilized in Algorithm 1. In this scenario, access to analytical posterior predictive distributions is unavailable, making the computation of common acquisition functions challenging. To address this, [2] propose a tailored acquisition function: a single posterior sample of $\alpha$ is produced, and the acquisition function is defined as $f_{\alpha}(\boldsymbol{x})-\lambda\|\boldsymbol{x}\|_{1}$, where a penalty term controlled by $\lambda$ is included. This acquisition function now has a closed-form, appearing as a quadratic form. Its optimization can be reformulated as a semidefinite program that can be efficiently approximated in polynomial time to a desired precision. With SBBO, there is no need for tailored acquisition functions as we just require having sampling access to the posterior predictive distribution. This enables using model (3.2) with the common acquisition functions in the BO literature.
NGBoost
@@
Natural Gradient Boosting (NGBoost, [8]) is a non-Bayesian algorithm for generic probabilistic prediction via gradient boosting. In NGBoost, the conditional distribution of the output given covariates is modeled through a parametric distribution $f(\boldsymbol{x})|\boldsymbol{x}\sim P_{\theta}(\boldsymbol{x})$. The parameters $\theta$ are obtained through an additive combination of $M$ base learners $b^{(m)}$ and an initial $\theta^{(0)}$,
$\theta=\theta^{(0)}-\eta\sum_{m=1}^{M}\rho^{(m)}\cdot b^{(m)}(x)$
@@
where $\rho^{(m)}$ are scaling factors and $\eta$ is a common learning rate. The base learners are trained to minimize a proper scoring rule using a refinement of gradient boosting. For this training process, NGBoost employs the natural gradient, a variant of gradient descent that takes into account the underlying geometry of probability distributions leading to faster convergence and improved stability.
@@
A notable strength of NGBoost lies in its adaptability, as it can be paired with various base learners. Our experiments focus on two such learners: shallow decision trees (referred to as NGBoost-dec) and linear regressions with lasso regularization (named NGBoost-linCV). For both setups, we use the log-score as the scoring rule, with the distribution $P_{\theta}$ set to a normal distribution in all experiments.
Bayesian Neural Network
@@
Bayesian Neural Networks (BNNs) are probabilistic deep learning models that replace deterministic weight values of a neural network with weight distributions. This allows the model to capture the existing epistemic uncertainty in the predictions due to limited training data. In our experiments, we employ a simple three hidden layer fully-connected BNN whose inputs are the one-hot encoded covariates. We train this model using variational inference, combined with the reparametrization estimator technique [17].
@@ 4 Experiments
We conduct experiments in three different combinatorial optimization problems [^1] All code to reproduce the experiments in this paper has been open sourced and is available at https://github.com/roinaveiro/sbbo.: a binary quadratic program with 10 variables, a contamination control in a food supply chain with 25 stages, and RNA sequence optimization with a sequence length of 30. We evaluate the performance of SBBO with the four probabilistic surrogate models from Section 3.3: Tanimoto Gaussian Process Regression (sbbo-GPr), Sparse Bayesian linear regression with interactions (sbbo-BLr), natural gradient boosting with a shallow decision tree model as base learner (sbbo-NGBdec), natural gradient boosting with an sparse linear regression as base learner (sbbo-NGBlinCV) and Bayesian Neural Network (sbbo-BNN). We compare the performance of these SBBO approaches with two benchmarks: simulated annealing (SA), a widely used algorithm in combinatorial optimization problems [26], and random local search (RS, [3]).
The experimental setup was as follows. We start with a random initial dataset containing five pairs of covariates and responses. For each probabilistic surrogate model and fixed number of iterations $t$, we follow these steps:
@@
    Use the surrogate model to determine the next evaluation point, employing the expected improvement as the acquisition function. Thus, the utility function is $u(\boldsymbol{x})=\max\left[f(\boldsymbol{x})-f^{*}(\boldsymbol{x}),0\right]$, where $f^{*}(\boldsymbol{x})$ represents the current estimate of the optimal true objective. As SBBO algorithm, we employ a Metropolis-Hastings sampling scheme, as in Algorithm 1. The proposal distribution randomly selects a covariate and assigns it a new value, uniformly sampled from its possible range. The chosen cooling schedule incrementally increases $H$ from 1 to 10000 in steps of 250.
@@
For each algorithm, we report the best function value obtained after $t$ evaluations of the true objective, averaged over 10 runs with the corresponding one standard error interval. The value of $t$ is chosen based on the specific problem setting.
@@ 4.1 Binary Quadratic Problem
Table 1: Average objective function values and margin of error for each algorithm after 120 iterations for the binary quadratic problem, 500 iterations for the contamination control problem and 300 iterations for the RNA problem. Bold values denote best value for each problem.
@@
The aim here is to maximize a quadratic function incorporating $l_{1}$ regularization. The objective function is expressed as $\boldsymbol{x}^{\top}Q\boldsymbol{x}-\lambda\|\boldsymbol{x}\|_{1}$, where $\boldsymbol{x}$ is a binary vector of dimension $d=10$. The matrix $Q$ is a random $d\times d$ matrix with entries generated as follows: each element is simulated as independent standard Gaussian with the resulting matrix then element-wise multiplied by another $d\times d$ matrix $K$, whose elements are $K_{ij}=\exp\left[-(i-j)^{2}/L_{c}^{2}\right]$, with the correlation length $L_{c}^{2}$ governing how rapidly the values of $K$ decay away from the diagonal. A larger value of $L_{c}^{2}$ results in a denser matrix $Q$, complicating the optimization task. In our experiments, we set $\lambda$ to 0 and fixed $L_{c}^{2}$ at 10.
Figure 1: Binary quadratic problem: average $\pm$ one standard error of objective function values by number of function evaluations. Large objective values indicate better performance.
The performance comparison of the different algorithms over 120 iterations is illustrated in Figure 1 and detailed in the second column of Table 1. As can be seen, SBBO outperforms RS and SA for most surrogate models. Notably, SBBO with sparse Bayesian linear regression achieves the highest objective value, closely followed by SBBO with NGB employing a sparse linear model as the base learner. An essential aspect is the assessment of performance with a limited number of objective function evaluations, as evaluating the true objective is the common bottleneck in most BO applications. Remarkably, SBBO-BLr, SBBO-BNN, and SA demonstrate good performance for a low number of evaluations ($<$25) (Fig. 1). However, it is worth noting that SA appears to get stuck in local optima around the 65th evaluation and SBBO-BNN performance does not grow as much as that of SBBO-BLr.
In this experiment, with a search space cardinality of 1024, it is feasible to compute the global optimum (11.24) via complete enumeration. Table 2 displays the deviation between the optimal solution discovered within 120 function evaluations for each algorithm and the global optimum. Notably, sbbo-BLr consistently identifies the optimal solution within 120 $\ll$ 1024 iterations across all 10 repetitions of the experiment.
Table 2: Binary quadratic problem: average distance to global optimum after 120 iterations with margin of error. Bold indicates best average performance.